OpenStudy (anonymous):

Sagari is making a cylindrical pottery container to hold beverages. She wants the container to hold at least one quart of liquad. (1 qt.= 57 3/4 cu. in., pie=22/7) If the inside diameter is three inches, how high must the inside height be. @smileyxl3

OpenStudy (smileyxl3):

\[57\frac{3}{4}=\frac{22}{7}(\frac{3}{2})^2h\] Solve for h. @cupcakestoner

OpenStudy (anonymous):

how do i solve for h?

OpenStudy (anonymous):

like i know what you mean, i just dont understand how to do the fractions and stuff.

OpenStudy (smileyxl3):

Do you have an answer choice of 8 1/6in.?

OpenStudy (anonymous):

ya, thanks:)